OpenStudy (anonymous):

4x^2 + ___x + 49 fill in the missing term of the perfect square trinomial

OpenStudy (anonymous):

28

OpenStudy (anonymous):

how'd you get 28?

OpenStudy (anonymous):

(2x+?)(2x+?)=4x^2 + ___x + 49 ?*? =49

OpenStudy (anonymous):

?=7 (2x+7)(2x+7) = 4x^2 + 28x + 49